A student pushes a desk 1.3 m across the floor at constant velocity. What is the mass of the desk if the coefficient of friction is 0.19 and the student did 49.4 J of work on the desk? 

 

 

 

 



1
Expert's answer
2021-06-03T09:19:10-0400

Let's first find the pushing force:


"W=F_{p}d,""F_p=\dfrac{W}{d}=\dfrac{49.4\ J}{1.3\ m}=38\ N."

Since the desk moves across the floor at constant velocity there is no acceleration:


"F_p-F_{fr}=0,""F_p=F_{fr}=\mu mg,""m=\dfrac{F_p}{\mu g}=\dfrac{38\ N}{0.19\cdot9.8\ \dfrac{m}{s^2}}=20.4\ kg."